Convert to the following decimal fractions to the form a\b where a and b€Z`b



1.3.1 0.1250



1.3.2 0.5




1.3.3 1.25



1.3.4 0.31


1
Expert's answer
2022-03-02T12:51:16-0500

1.3.1


"0.125=\dfrac{125}{1000}=\dfrac{1}{8}"

1.3.2


"0.5=\dfrac{5}{10}=\dfrac{1}{2}"

1.3.3


"1.25=\dfrac{125}{100}=\dfrac{5}{4}"

1.3.4


"0.31=\dfrac{31}{100}"